Harris Company manufactures and sells a single product. A partially completed schedule of the company’s total costs and costs per unit over the relevant range of 51,000 to 91,000 units is given below:

Units Produced and Sold
51,000 71,000 91,000
Total costs:
Variable costs $112,200
Fixed costs 380,000
Total costs $492,200 $0 $0
Cost per unit:
Variable cost
Fixed cost
Total cost per unit $0.00 $0.00 $0.00


Required:

1. Complete the schedule of the company’s total costs and costs per unit as given in the relevant tab below.

2. Assume that the company produces and sells 81,000 units during the year at a selling price of $8.55 per unit. Prepare a contribution format income statement for the year.
Answer #1
  • Requirement1

Units Produced and Sold

51,000

71,000

91,000

Total costs:

Variable costs

$112,200

$156,200

$200,200

Fixed costs

$380,000

$380,000

$380,000

Total costs

$492,200

$536,200

$580,200

Cost per unit:

Variable cost

$2.20

$2.20

$2.20

Fixed cost

$7.45

$5.35

$4.18

Total cost per unit

$9.65

$7.55

$6.38

--Working

Units Produced and Sold

51000

71000

91000

Total costs:

Variable costs

112200

=71000*2.2

=91000*2.2

Fixed costs

380000

380000

380000

Total costs

=112200+380000

=156200+380000

=200200+380000

Cost per unit:

Variable cost

=112200/51000

2.2

2.2

Fixed cost

=380000/51000

=380000/71000

=380000/91000

Total cost per unit

=2.2+7.45

=2.2+5.35

=2.2+4.18

  • Contribution Income Statement

Sales

[81000 units x $ 8.55]

$692,550

Variable cost

[81000 units x $ 2.20]

$178,200

Contribution margin

$514,350

Fixed cost

$380,000

Net Income

$134,350